# CodeBrain: Bridging Decoupled Tokenizer and Multi-Scale Architecture for EEG Foundation Model

### v1

@@ 4.4 Ablation Study
Dataset Codebook SWA CL Cohen’s Kappa Weighted F1 Balanced Accuracy
FACED 9-Class Dual 0.5270 $\pm$ 0.0039 0.5849 $\pm$ 0.0033 0.5816 $\pm$ 0.0028
@@ 4.5 Vector Visualization
To better understand how the proposed TFDual-Tokenizer overcomes the limitations of single-codebook quantization in handling heterogeneous EEG signals, we visualize its learned temporal and frequency codes on the ISRUC_S3 dataset. As shown in Figure 4 (a) and (b), both temporal and frequency codebooks exhibit clustering behavior, capturing domain-specific features, such as characteristic waveforms or frequency patterns associated with specific sleep stages (e.g., N3). Figure 4 (c) and (d) further demonstrate that our decoupled design enables flexible composition: the same temporal code can pair with different frequency codes to represent different sleep stages, and vice versa. This indicates that temporal and frequency representations contribute both independently and complementarily to the final task representation. Such tokenizer supports a richer and more structured representation space for heterogeneous features, improving both performance and interpretability.
@@ 5 Conclusion
In this work, we propose CodeBrain, an efficient EEG foundation model that integrates three key components: a temporal–frequency decoupled tokenizer for heterogeneous representation learning, a structured state space backbone for modeling sparse global dependencies, and a sliding window attention mechanism for capturing local intra-patch patterns. Through comprehensive evaluation, CodeBrain consistently outperforms prior foundation models under linear probing, demonstrating strong generalization and transferability. Our results highlight the potential of CodeBrain as a versatile backbone for EEG-based brain–computer interface applications, offering a scalable and interpretable foundation for future developments in neural decoding and AI in neuroscience.
Appendix A Preliminaries
Convolution State Space Models
@@ Appendix A Preliminaries
where $u(t)$ is a vector that represents the input of the system, while $y(t)$ is a vector that represents the output of the system. $x(t)$ and its derivative $x^{\prime}(t)$ represent the latent states of the system, typically in the form of an N-D vector. And $A,B,C,D$ here are the state, input, output, and feedforward matrices, defining the relationship between the input, output, and state vector. Following Gu et al. [75], $\bar{D}$ is set equal to 0 since it can be replaced by the residual connection. Now, Equation 18 resembles an architecture similar to RNN, allowing us to recurrently compute $x_{k}$. Let the initial state be $x_{k-1}=0$, and we can unroll Equation 18 as follows:
@@
Appendix B Pretraining Results
@@
B.1 TFDual-Tokenizer Pretraining Results
@@
During pretraining, we track the number of unused codes in both the temporal and frequency codebooks of the TFDual-Tokenizer, each with a size of 4096. As shown in Figure 7, the frequency codebook demonstrates a rapid decrease in unused codes, while the temporal codebook shows a slower and more incremental reduction. A more detailed analysis of the temporal–frequency complementarity is provided in Section C.
@@
B.2 EEGSSSM Pretraining Results
@@
Appendix C Class-Specific Token Ratio Analysis for TFDual-Tokenizer
$\text{Dominance}(c)=\frac{\max_{y}N_{c}^{(y)}}{\sum_{y}N_{c}^{(y)}}$ (21)
[…]
Figure 9: Class-Specific Code Ratio Across Different Codebooks.
Appendix D Improving Temporal Codebook Learning via Contrastive Loss
@@
Appendix E Dataset Description
@@
Appendix F Baselines and Metrics Description
F.1 Metrics
@@ F.2 Baselines
Appendix G Hyperparameter Setting
@@ G.1 Pretraining Settings
Table 5: Hyperparameters for EEGSSM.
@@
G.2 Fine-tuning Settings on Downstream Tasks
@@
Appendix H Model Comparison
Backbone Efficiency Comparison
To evaluate the computational efficiency of our proposed SGConv module, we conduct an ablation study by replacing it with three common sequence modeling modules: CNN, LSTM, and Transformer. We compare their model sizes, floating-point operations (FLOPs), and iteration times, as shown in Figure 11. Specifically, the CNN variant uses a three-layer depthwise separable convolution block, while the LSTM and Transformer variants use a single layer of standard LSTM and Transformer Encoder (implemented by Pytorch), respectively. In terms of parameter count, SGConv contains only 8.74M parameters, significantly fewer than CNN (28.22M), LSTM (30.35M), and Transformer (27.79M). For FLOPs, SGConv also achieves the lowest computational cost at 15.17G, compared to Transformer’s 21.2G. Regarding iteration time, SGConv is slightly slower than Transformer and CNN models in terms of training speed, but it outperforms the LSTM model. In summary, SGConv effectively reduces the number of parameters while maintaining computational complexity, which helps the model to be trained and inferred on smaller GPUs.
Additional Evaluation on Other BCI Tasks
We report the performance of CodeBrain on four additional EEG datasets not included in the main text, covering diverse domains of sleep staging, motor imagery, event detection, and abnormality classification in Tables 7 to 10. These allow us to assess the cross-domain generalization ability of our pretrained model beyond the main text.
We note that both TUAB and TUEV originate from the TUH EEG corpus [66], which overlaps with our pretraining source (TUEG). To avoid overfitting to this distribution and promote generalization, we stop pretraining at epoch 10 as discussed in Section B.2. While this may limit gains on TUH datasets compared to previous EFM, such as CBraMod (trained for 40 epochs in the same pretraining dataset) [19], CodeBrain still achieves superior or competitive results.
@@ Appendix H Model Comparison
As shown in Table 9, CodeBrain achieves the highest Cohen’s Kappa (0.6912, an improvement of +0.0140 over the best baseline [19]) and Weighted F1 (0.8362) on TUEV. Although its Balanced Accuracy is lower than CBraMod [19], we attribute this to reduced sensitivity on the rare SPSW class. Since TUEV shares distributional overlap with our pretraining source (TUEG), we stop pretraining at epoch 10 to prevent overfitting, unlike CBraMod’s 40-epoch training as discussed in subsection B.2. We also report LaBraM’s results based on its original 23-channel setting [20], while CodeBrain follows the 16-channel configuration used in CBraMod. Similarly, EEGPT [18] does not adopt a linear fine-tuning protocol but applies two convolutional layers before entering the foundation model, followed by an MLP head. While such architectural choices may enhance performance, we follow their respective fine-tuning settings. In such cases, CodeBrain still outperforming both LaBraM and EEGPT under their own fine-tuning settings clearly demonstrates the robustness of our approach.
@@
As shown in Table 10, CodeBrain achieves the highest Balanced Accuracy (0.8294) on TUAB, slightly outperforming CBraMod [19]. Similar to TUEV, TUAB is part of the TUH EEG corpus family [66] and thus closely aligned with our pretraining source (TUEG). As discussed in subsection B.2, we adopt an early stopping strategy at epoch 10 to mitigate overfitting to this distribution, which may partly account for the slightly lower AUROC and AUC-PR compared to CBraMod, trained for 40 epochs on the same dataset. While LaBraM leverages a 23-channel montage [20] and EEGPT [18] employs two convolutional layers before the foundation model, we retain their respective fine-tuning protocols for comparison. Despite these potentially stronger configurations, CodeBrain still exceeds both models under their own settings, highlighting its strong and consistent generalization.
@@
Appendix I Ablation on Mask Ratio
@@
Appendix J Detailed Results on Scaling Laws
@@
We provide the detailed scaling law results for both data and model size across three representative EEG datasets (FACED, SEED-V, and ISRUC_S3), covering three evaluation metrics. For brevity, only Cohen’s kappa scores are included in the main text, while full results are provided in this section. Prior work [19, 20] has explored the effect of scaling EEG foundation models using 1 to 1000 hours of pretraining data and model sizes up to 5M parameters. We extend this analysis in two key dimensions:

### v2

@@ 4.3 Comparison with Baselines
Dataset Codebook CL SWA SGConv Gate Cohen’s Kappa Weighted F1 Balanced Accuracy
FACED 9-Class Dual 0.5406 $\pm$ 0.0084 0.5953 $\pm$ 0.0113 0.5941 $\pm$ 0.0098
@@
Dual 0.6218 $\pm$ 0.0427 0.6956 $\pm$ 0.0279 0.6664 $\pm$ 0.0316
Dual 0.5478 $\pm$ 0.0302 0.6429 $\pm$ 0.0307 0.6258 $\pm$ 0.0448
4.4 Ablation Study
We conduct ablation studies on three datasets with the same five seeds as in the main experiments to evaluate the key components of CodeBrain (Table 3). Below is the detailed analysis:
(1) Tokenizer configuration: we compare the proposed TFDual-Tokenizer (Dual) with variants using a single domain codebook (Temporal or Frequency), or a shared codebook jointly reconstructing both domains (Mixed). Across all datasets, the Dual codebook consistently yields superior performance.
(2) Contrastive learning (CL): We evaluate the impact of contrastive learning in TFDual-Tokenizer pretraining. It leads to consistent gains, indicating a better capture of temporal patterns. Moreover, the CL facilitates convergence of the temporal codebook, with detailed analyses provided in Appendix E.
(3) Components of EEGSSM: Evaluating the SWA, SGConv, and Gate modules, which demonstrate improvements in EEG representation learning.
(4) Scaling Laws: Prior works (wang2024cbramod, ; jianglarge, ) explored scaling with 1-1000 hours of EEG data for pretraining. We extend to 1k-9k hours and 3M-150M models. As shown in Figure 3, Kappa consistently improves with more data and parameters. These results confirm that larger models yield consistent but diminishing returns. More detailed results are in Appendix M.
We further investigate several key design choices, including mask ratio, SWA window size, codebook size, and patch size in Appendix J. We report robustness experiments to demonstrate the adaptability of dynamic positional embeddings in Appendix L, and computational efficiency in Appendix K. We also provide low-resource comparisons (Appendix N) and pretraining curves (Appendix D), showing stable VQ convergence and efficient EEGSSM training.
@@ 4.5 Vector Visualization
To demonstrate how the TFDual-Tokenizer models heterogeneous EEG, we visualize its learned temporal and frequency codes on the ISRUC_S3 dataset. As shown in Figure 4(a)(b), each domain-specific codebook captures meaningful structures: temporal codes correspond to neural events (e.g., slow waves), while frequency codes highlight spectral rhythms such as dominant delta activity, either of which can support sleep stage discrimination. Yet in many cases a single domain is insufficient, and discrimination arises only from their composition, as in Figure 4(c)(d), where the same temporal code can pair with different frequency codes, and vice versa, to yield complementary representations. This decoupled design expands the representation space and enhances interpretability, with additional potential links to neurophysiological features and quantitative analyses provided in Appendix A.
@@ 5 Conclusion
In this paper, we present CodeBrain, a EEG foundation model that unifies interpretable tokenization with a brain-inspired multi-scale architecture. The TFDual-Tokenizer decouples heterogeneous EEG signals, expanding the representation space while suggesting domain-specific interpretability, and the EEGSSM architecture integrates structured global convolution with sliding-window attention to efficiently capture both long-range and local dependencies. Pretrained on the large-scale TUEG corpus, CodeBrain demonstrates strong generalization across 8 tasks and 10 datasets under distribution shifts, with comprehensive ablations, scaling-law analyses confirming robustness and scalability. These results establish CodeBrain as a strong foundation for neural time-series representation learning.
Appendix A Interpretability Analyses of TFDual-Tokenizer Representations
We conduct interpretability analyses of the decoupled TFDual-Tokenizer. As case studies, we visualize selected tokens in relation to well-established physiological markers, showing that frequency codes tend to align with spectral rhythms, whereas temporal codes correspond to neural events. These qualitative examples highlight that the tokenizer does not discretize signals in an abstract manner, but captures domain-specific structures with clear clinical relevance. Finally, we complement these visualizations with a quantitative analysis of class-specific token usage on 4 datasets, demonstrating that the learned tokens are not only interpretable but also discriminative, contributing to improved downstream classification.
A.1 Exploring Frequency Token Patterns in Relation to Spectral Rhythms
To illustrate the interpretability of the learned frequency tokens, we take sleep staging as an example downstream task and analyze token embeddings. We focus on N2 and N3 stages, since they are characterized by the most distinctive spectral rhythms in clinical sleep scoring: spindle in the sigma band (11-16 Hz) for N2, and slow wave in the delta band (0.5-4 Hz) for N3.
As shown in Figure 5, frequency code No. 2298 captures a sigma bump, with prominent peaks localized in the spindle range. This aligns with the definition of N2 [68], which identifies spindle as a key marker. Quantitatively, this code exhibited a clear sigma peak in approximately $15.4\%$ of its assigned patches. Similarly, Figure 6 shows that frequency code No. 32 predominantly encodes delta dominance, a typical spectral feature of N3 sleep. This code displayed clear delta dominance in about $18.3\%$ of its assigned segments. Taken together, these findings suggest that our frequency-branch tokenizer helps establish a frequency vocabulary for EEG, suggesting potential links between the learned codes and clinically interpretable oscillations beyond class frequency bias.
Figure 5: Frequency code capturing sigma bump, a typical spectral rhythm of N2 stage.
Figure 6: Frequency code capturing delta dominance, a typical spectral rhythm of N3 stage.
A.2 Exploring Temporal Token Patterns in Relation to Neural Events
In addition to spectral rhythms, we further analyzed temporal codes on the sleep staging dataset to assess whether the tokenizer captures characteristic neural events in raw EEG. Figure 8 and Figure 7 show that certain codes are closely associated with K-complexes and sleep spindles, the hallmark waveforms of N2 sleep. Quantitatively, $\approx 5.34\%$ of the assigned patches contained K-complexes and $\approx 7.01\%$ contained spindles. These rates are consistent with clinical expectations (1–2 K-complexes and 2–5 spindles per minute in N2 [68]) and, when considering the overall prevalence of N2 in five-class sleep staging, still substantially exceed what would be expected from random token assignment.
Figure 7: Temporal code capturing sleep spindle, a typical neural events of N2 stage.
Figure 8: Temporal code capturing K complex, a typical neural events of N2 stage.
Figure 9 shows that temporal code No. 1537 corresponds to slow waves, the defining feature of N3 sleep, occurring in $\approx 40.4\%$ of its assigned patches. This exceeds the $20\%$ per-epoch criterion for N3 scoring [68] and, given the overall prevalence of N3, indicates that this token provides a meaningful link to neural events.
Figure 9: Temporal code capturing slow wave, a typical neural events of N3 stage.
Taken together, these observations suggest that the temporal branch of our tokenizer contributes to establishing a vocabulary of neurophysiological events, complementing the frequency-domain findings. While temporal waveforms are often noisier and harder to model than spectral rhythms [16], our decoupled design allows temporal tokens to emerge with meaningful associations to critical neural events. This indicates that the temporal branch offers useful links to clinically relevant waveforms and demonstrates the effectiveness of our approach in capturing complementary structure.
A.3 Class-Specific Token Ratio Analysis for TFDual-Tokenizer
$\text{Dominance}(c)=\frac{\max_{y}N_{c}^{(y)}}{\sum_{y}N_{c}^{(y)}}$ (17)
$\text{Class-Specific Token Ratio}=\frac{\text{\# class-specific tokens}}{\text{Total tokens in codebook}}$ (18)
Figure 10: Class-Specific Code Ratio Across Different Codebooks.
Appendix B Structure State Space Model
@@
where $u(t)$ is a vector that represents the input of the system, while $y(t)$ is a vector that represents the output of the system. $x(t)$ and its derivative $x^{\prime}(t)$ represent the latent states of the system, typically in the form of an N-D vector. And $A,B,C,D$ here are the state, input, output, and feedforward matrices, defining the relationship between the input, output, and state vector. Following Gu et al. [69], $\bar{D}$ is set equal to 0 since it can be replaced by the residual connection. Now, Equation 20 resembles an architecture similar to RNN, allowing us to recurrently compute $x_{k}$. Let the initial state be $x_{k-1}=0$, and we can unroll Eq 20 as follows:
@@
While Eq. 22 shows that the SSM can be computed using FFT-based convolution, it is important to formalize the complexity guarantee. We now state the following proposition.
Proposition 1. Let $(A,B,C,D)$ denote the discretized state-space matrices of an S4 layer, with input sequence $u\in\mathbb{R}^{N}$. The output $y\in\mathbb{R}^{N}$ can be written as a linear convolution y=kuy=k*u with kernel
$k_{0}=D,\quad k_{n}=CA^{n-1}B,\quad n\geq 1.$
Using the convolution theorem and the FFT, this convolution can be computed in time $\Theta(N\log N)$.
Proof. This proof follows the proof in Lemma C.2 of gu2021efficiently [49]:
Expanding the recurrence
$x_{n+1}=Ax_{n}+Bu_{n},\qquad y_{n}=Cx_{n}+Du_{n},$ (23)
yields
$y_{n}=Du_{n}+\sum_{t=0}^{n-1}CA^{\,n-1-t}B\,u_{t}=\sum_{t=0}^{n}k_{n-t}\,u_{t},$ (24)
where $k_{0}=D$ and $k_{n}=CA^{n-1}B$. Thus y=kuy=k*u.
Let $\mathcal{F}_{N}$ denote the $N$-point DFT. By the convolution theorem,
$\mathcal{F}_{N}(y)=\mathcal{F}_{N}(k)\odot\mathcal{F}_{N}(u),$ (25)
so that
$y=\mathcal{F}_{N}^{-1}\!\left(\mathcal{F}_{N}(k)\odot\mathcal{F}_{N}(u)\right).$ (26)
The cost of forward FFT, element-wise product, and inverse FFT is $\Theta(N\log N)$.
The S4 parameterization ensures $A$ admits a diagonal-plus-low-rank (DPLR) structure, so the kernel takes the form
$k_{n}=\sum_{j=1}^{r}\alpha_{j}\,\lambda_{j}^{\,n-1},\qquad r\ll N.$ (27)
Each exponential sequence can be generated recursively in $\mathcal{O}(N)$, yielding all $k_{0},\ldots,k_{N-1}$ in linear time.
Combining the above,
$T(N)=\underbrace{\Theta(N\log N)}_{\text{FFT convolution}}+\underbrace{\mathcal{O}(N)}_{\text{kernel generation}}=\Theta(N\log N).$ (28)
Appendix C Theoretical Analysis of Decoupled Codebook Training
Proposition 2.1. Let $X=(X_{t},X_{f})$ denote the temporal and frequency representations of an EEG segment, assumed approximately independent. Consider an additive reconstruction distortion $d(x,\hat{x})=d_{t}(x_{t},\hat{x}_{t})+d_{f}(x_{f},\hat{x}_{f})$. Under a fixed total codebook size $K=2^{R}$, a product codebook $\mathcal{C}_{t}\times\mathcal{C}_{f}$ with $|\mathcal{C}_{t}|=2^{R_{t}}$ and $|\mathcal{C}_{f}|=2^{R_{f}}$, $R_{t}+R_{f}=R$, achieves a minimum expected distortion
$D_{\text{prod}}^{\star}(R)=\min_{R_{t}+R_{f}=R}\Bigl(D_{t}^{\star}(R_{t})+D_{f}^{\star}(R_{f})\Bigr),$ (29)
which satisfies
$D_{\text{mix}}^{\star}(R)\;\geq\;D_{\text{prod}}^{\star}(R),$ (30)
where $D_{\text{mix}}^{\star}(R)$ is the minimum distortion of a single mixed codebook $\mathcal{C}_{\text{mix}}\subset\mathbb{R}^{d_{t}+d_{f}}$ of size $2^{R}$.
Proof. The argument combines the separability of the rate–distortion (R–D) function for independent sources under additive distortion and the high-rate quantization approximation to the Shannon R–D limit.
Let $X=(X_{t},X_{f})$ with $X_{t}\perp X_{f}$, and distortion
$d(x,\hat{x})\;=\;d_{t}(x_{t},\hat{x}_{t})+d_{f}(x_{f},\hat{x}_{f}).$ (31)
Then the Shannon R–D function satisfies
$R(D)\;=\;\min_{D_{t}+D_{f}=D}\,\Bigl(R_{t}(D_{t})+R_{f}(D_{f})\Bigr),$ (32)
where $R_{t}(\cdot)$ and $R_{f}(\cdot)$ are the marginal R–D functions for $X_{t}$ and $X_{f}$. By convex duality, the optimal test channel factorizes:
$p(\hat{x}_{t},\hat{x}_{f}\mid x_{t},x_{f})\;=\;p(\hat{x}_{t}\mid x_{t})\,p(\hat{x}_{f}\mid x_{f}),$ (33)
and the optimal distortion allocation solves
$\min_{D_{t},D_{f}}\;R_{t}(D_{t})+R_{f}(D_{f})\quad\text{s.t. }D_{t}+D_{f}=D,$ (34)
with KKT condition $R_{t}^{\prime}(D_{t}^{\star})=R_{f}^{\prime}(D_{f}^{\star})$.
At rate $R_{t}$ bits for $X_{t}$ and $R_{f}$ bits for $X_{f}$ (with $R_{t}+R_{f}=R$), the minimal achievable distortions are $D_{t}^{\star}(R_{t})$ and $D_{f}^{\star}(R_{f})$. In the high-rate regime, practical vector quantizers approach these Shannon limits, yielding
$D_{\text{prod}}(R_{t},R_{f})\;\approx\;D_{t}^{\star}(R_{t})+D_{f}^{\star}(R_{f}).$ (35)
Optimizing over all feasible splits gives
$D_{\text{prod}}^{\star}(R)\;=\;\min_{R_{t}+R_{f}=R}\bigl(D_{t}^{\star}(R_{t})+D_{f}^{\star}(R_{f})\bigr).$ (36)
Any mixed codebook $\mathcal{C}_{\text{mix}}\subset\mathbb{R}^{d_{t}+d_{f}}$ with $|\mathcal{C}_{\text{mix}}|=2^{R}$ cannot beat the Shannon R–D limit, so
$D_{\text{mix}}^{\star}(R)\;\geq\;D^{\star}(R).$ (37)
But from (34)– (35),
$D^{\star}(R)\;=\;\min_{R_{t}+R_{f}=R}\bigl(D_{t}^{\star}(R_{t})+D_{f}^{\star}(R_{f})\bigr)\;=\;D_{\text{prod}}^{\star}(R).$ (38)
Therefore,
$D_{\text{mix}}^{\star}(R)\;\geq\;D_{\text{prod}}^{\star}(R),$ (39)
Appendix D Pretraining Results
@@
D.1 TFDual-Tokenizer Pretraining Results
@@
During pretraining, we track the number of unused codes in both the temporal and frequency codebooks of the TFDual-Tokenizer, each with a size of 4096. As shown in Figure 13, the frequency codebook demonstrates a rapid decrease in unused codes, while the temporal codebook shows a slower and more incremental reduction. A more detailed analysis of the temporal-frequency complementarity is provided in Section A.3.
@@
D.2 EEGSSSM Pretraining Results
@@
Appendix E Improving Temporal Codebook Learning via Contrastive Loss
@@
Appendix F Dataset Description
@@
Appendix G Baselines and Metrics Description
G.1 Metrics
@@ G.2 Baselines
Appendix H Hyperparameter Setting
@@ H.1 Pretraining Settings
Table 5: Hyperparameters of Pre-training.
@@
H.2 Parameters of EEGSSM
The model architecture parameters used in EEGSSM during pre-training are shown in Table 6.
Table 6: Configuration of EEGSSM
Parameters Values
Input size 200
Hidden dimension 200
Output size 200
Number of layers 8
Max sequence length 570
SGConv state 64
SGConv bidirectional True
Layer normalization True
Sliding window attention length 1s
H.3 Fine-tuning Settings on Downstream Tasks
@@
Appendix I Additional Evaluation on Other BCI Tasks
We report the performance of CodeBrain on four additional EEG datasets not included in the main text, covering diverse domains of sleep staging, motor imagery, event detection, and abnormality classification in Tables 8 to 11. These allow us to assess the cross-domain generalization ability of our pretrained model beyond the main text.
We note that both TUAB and TUEV originate from the TUH EEG corpus [33], which overlaps with our pretraining source (TUEG). To avoid overfitting to this distribution and promote generalization, we stop pretraining at epoch 10 as discussed in Section D.2. While this may limit gains on TUH datasets compared to previous EFM, such as CBraMod (trained for 40 epochs in the same pretraining dataset) [15], CodeBrain still achieves superior or competitive results.
@@ Appendix I Additional Evaluation on Other BCI Tasks
As shown in Table 10, CodeBrain achieves the highest Cohen’s Kappa (0.6912, an improvement of +0.0140 over the best baseline [15]) and Weighted F1 (0.8362) on TUEV. Although its Balanced Accuracy is lower than CBraMod [15], we attribute this to reduced sensitivity on the rare SPSW class. Since TUEV shares distributional overlap with our pretraining source (TUEG), we stop pretraining at epoch 10 to prevent overfitting, unlike CBraMod’s 40-epoch training as discussed in subsection D.2. We also report LaBraM’s results based on its original 23-channel setting [16], while CodeBrain follows the 16-channel configuration used in CBraMod. Similarly, EEGPT [14] does not adopt a linear fine-tuning protocol but applies two convolutional layers before entering the foundation model, followed by an MLP head. While such architectural choices may enhance performance, we follow their respective fine-tuning settings. In addition, following the experimental setup of CBraMod, we also report the results of experiments conducted by removing all TUEV and TUAB samples from the TUEG dataset. It can be seen that although our model’s performance slightly declined on TUEV, it still surpassed other baselines. In such cases, CodeBrain still outperforming both LaBraM and EEGPT under their own fine-tuning settings clearly demonstrates the robustness of our approach.
@@
As shown in Table 11, CodeBrain achieves the highest Balanced Accuracy (0.8294) on TUAB, slightly outperforming CBraMod [15]. Similar to TUEV, TUAB is part of the TUH EEG corpus family [33] and thus closely aligned with our pretraining source (TUEG). As discussed in subsection D.2, we adopt an early stopping strategy at epoch 10 to mitigate overfitting to this distribution, which may partly account for the slightly lower AUROC and AUC-PR compared to CBraMod, trained for 40 epochs on the same dataset. While LaBraM leverages a 23-channel montage [16] and EEGPT [14] employs two convolutional layers before the foundation model, we retain their respective fine-tuning protocols for comparison. In addition, the impact of duplicate data in the pre-training set on our model is smaller on the TUAB dataset, and in some experiments it can even surpass situations without leaking. This may be due to the larger size of the TUAB dataset. Despite these potentially stronger configurations, CodeBrain still exceeds both models under their own settings, highlighting its strong and consistent generalization.
@@
Appendix J Ablation on Design Choices
J.1 Ablation on Mask Ratio
@@
J.2 Ablation on SWA Window Size
We conduct an ablation study to investigate the effect of the SWA window size in the EEGSSM framework. The window size of SWA means the length of the segment observed by the attention mechanism in the SWA mechanism. Usually, the window size of SWA is an odd number because the model generally focuses not only on a single segment but also needs to observe features in adjacent segments. Therefore, its size is usually $2n+1$, where $n$ represents the length of SWA. As shown in Figure 18 and Figure 19, SWA window size = 1 achieved the best performance on both datasets among the 5 SWA window sizes. With the increase of the SWA window size, the SWA is close to self-attention. Therefore, as the SWA window size increases, the model’s performance actually improves because more data will be involved in the attention calculation. When the SWA window size is 1, it is a special case where SWA only calculates for one second. This is equivalent to the model performing attention mechanism calculations within one second. Overall, SWA can help the model achieve a certain performance improvement.
Figure 18: Performance of different SWA window sizes on the FACED dataset.
Figure 19: Performance of different SWA window sizes on the SEED-V dataset.
J.3 Ablation on Codebook Size
The size of the codebook is an important parameter; a codebook that is too large may lead to unstable training, while a codebook size that is too small may result in mixed information. Ideally, the codebook should maintain a small amount of unused codes but not be 0. We tested several combinations of different time-domain and frequency-domain codebook sizes to observe their unused codes during Tokenizer training.
Table 15 shows the unused temporal codes and unused frequency codes under different codebook size combinations. When both the Temporal codebook size and Frequency codebook size are set to 2048, the unused frequency code is 0, indicating that there are duplicate frequency codes in the current codebook. When choosing a codebook size of 8192, the unused temporal codes reached 3401 and the unused frequency codes reached 1260, indicating that a large number of codes in the completed training codebook were not used. We ultimately selected 4096 for both. Although the frequency domain often yields richer representations, enlarging its codebook may increase reliance on it, so we kept temporal and frequency codebooks equal. Considering capacity and utilization, 4096–4096 offers the best trade-off.
Table 15: Temporal and frequency codebook statistics
Temporal Codebook Size Frequency Codebook Size Unused Temporal Codes Unused Frequency Codes
2048 2048 12 0
2048 4096 116 0
4096 4096 225 165
4096 8192 321 931
8192 8192 3401 1260
J.4 Ablation on Patch Size
The size of the patch window is also an important adjustable parameter, which affects temporal resolutions and masking strategies. To explore the impact of patch window size on the model, we used window sizes ranging from 0.5s to 5s for complete two-stage pre-training and full-parameter fine-tuning. Note that for some datasets where the patch size is larger than the channel length, such as SEED-V, we pad the portion exceeding the available data to match the patch size in this experiment.
Table 16 and Table 17 show the performance of our method with different patch sizes on the SEED-V dataset and ISRUC_3 dataset. Notably, in SEED‑V, patch lengths longer than 1s require heavy padding, causing large performance drops; in ISRUC_S3, the shortest 0.5s patches achieve the worst performance, likely because they fragment key waveforms in sleep staging (e.g., Spindle $\geq$0.5s). From these results, the 1s setting is supported by two key considerations:
Broad compatibility with downstream task. 1s is a divisor of most downstream sequence lengths (1–30s), minimizing padding and ensuring transferability. For example, on the SEED-V dataset, if the patch size chosen by the model is greater than 1s, some methods (such as padding) need to be adopted to enable model training. These methods may usually impair the model’s performance because they introduce additional noise or increase computational load. EEG datasets with durations less than 1 second are relatively rare, as most datasets have at least 1 second of data. If the data does not exactly match the whole seconds, the cost of processing such a dataset is also relatively small. Prior EEG foundation models (e.g., LaBram [16], CBraMod [15]) also adopt 1s patches for this reason.
Semantic integrity. Choosing a 1s patch length preserves the natural structure of EEG waveforms and prevents semantic fragmentation. Many physiologically meaningful EEG events have characteristic durations: for example, K-complexes are typically around 1s, spindles last 0.5–2s, and event-related potentials such as P300 occur in the range of 0.3–0.6s. If patches are shorter than these characteristic scales, the temporal branch of the TFDual-Tokenizer may only capture partial fragments of these waveforms, leading to loss of semantic context.
Table 16: Performance on SEED-V with different patch sizes
Patch Length Cohen’s Kappa Weighted F1 Balanced Accuracy
[…]
Table 17: Performance on ISRUC_3 with different patch sizes
Patch Length Cohen’s Kappa Weighted F1 Balanced Accuracy
[…]
Appendix K Backbone Efficiency Comparison
To evaluate the computational efficiency of our proposed SGConv module, we conduct an ablation study by replacing it with three common sequence modeling modules: CNN, LSTM, and Transformer in EEGSSM block. We compare their model sizes, floating-point operations (FLOPs), and iteration times, as shown in Figure 20. Specifically, the CNN variant uses a 3-layer depthwise separable convolution block, while the LSTM and Transformer variants use a single layer of standard LSTM and Transformer Encoder (implemented by Pytorch), respectively. In terms of parameter count, SGConv contains 15.17M parameters, fewer than CNN (16.22M), LSTM (17.35M), and Transformer (21.2M). For FLOPs, SGConv also achieves the lowest computational cost at 8.74G, compared to Transformer’s 27.79G. Regarding iteration time, SGConv is slightly slower than Transformer and CNN models in terms of training speed, but it outperforms the LSTM model. In summary, SGConv effectively reduces the number of parameters while maintaining computational complexity, which helps the model to be trained and inferred on smaller GPUs.
Appendix L Channel Robustness
In real-world scenarios, the collection of EEG often encounters situations where channels are missing, especially when using machines from different manufacturers. To test the model’s performance on datasets with missing channel data, we randomly mask some channels in the training data for full parameter fine-tuning. We selected the FACED and SEED-V datasets for experiments because they represent short-sequence and long-sequence cases respectively, and their numbers of channels are relatively complete.
Figure 21: Performance after randomly masking different numbers of channels during the full parameter fine-tuning stage.
We evaluate the performance of the CodeBrain model and CBraMod model in scenarios with missing channels. We conducted three different experiments, randomly masking 12.5%, 25%, and 50% of the channels in each experiment, respectively. Figure 21 shows results of the experiment. It can be seen that our CodeBrain model outperforms the CBraMod model in all channel masking scenarios. On the FACED dataset, our model’s performance after masking 25% of the channels is still close to that of CBraMod without masking. The performance decline of the CBraMod model is also faster than that of our model. This trend is even more pronounced on the SEED-V dataset. Through this experiment, we can demonstrate that the channel robustness of CodeBrain is stronger than that of the CBraMod model, retaining most of its performance even in cases of channel failure.
Appendix M Detailed Results on Scaling Laws
@@
We provide the detailed scaling law results for both data and model size across three representative EEG datasets (FACED, SEED-V, and ISRUC_S3), covering three evaluation metrics. For brevity, only Cohen’s kappa scores are included in the main text, while full results are provided in this section. Prior work [15, 16] has explored the effect of scaling EEG foundation models using 1 to 1000 hours of pretraining data. We extend this analysis in two key dimensions: